# Unveiling Visual Perception in Language Models: An Attention Head Analysis Approach

### v1

@@ 1 Introduction
The integration of large language models (LLMs) into multimodal tasks has led to impressive advancements, with MLLMs, or large vision-language models (LVLMs), achieving notable success in areas ranging from traditional image captioning to more complex visual dialogues. A key element in training these models is the adapter, which connects the visual encoder and the LLM. The standard approach typically begins by freezing both the visual encoder (often based on models like CLIP [28]) and the LLM, while fine-tuning only the adapter to map visual embeddings into a language-compatible space. Subsequently, the LLM is fine-tuned to follow instructions for visual tasks, enabling it to become a versatile model capable of solving vision problems that traditionally required multiple specialized models. This approach works remarkably well, raising an intriguing question: does the adapter merely translate visual embeddings into a language space that the LLM can interpret?
@@
Recent studies [14, 36] reveal that representations from encoders like CLIP can be effectively transformed into a space compatible with LLMs by training an adapter, whereas other visual representations face challenges in achieving similar results. This raises a fundamental question about how LLMs, trained exclusively on language embeddings, interact with embeddings originating from fundamentally different visual spaces. In particular, Chan et al. [5] emphasizes that the embedding spaces for visual tokens differ significantly from those used for language, as visual tokens do not function as direct counterparts to language tokens but instead occupy a distinct space. Therefore, an in-depth analysis of how LLMs process embeddings from such a divergent space is essential, as current research lacks a clear understanding of how LLMs attend to these visual tokens during inference. To address this problem, it is necessary to explore the attention mechanisms of LLMs when handling visual tokens.
@@
In parallel, research within the LLM domain has increasingly focused on uncovering the internal mechanisms that drive LLM capabilities, particularly around specialized attention heads that contribute distinct functions. For example, “Duplicate Heads” enhance the model’s ability to focus on repeated tokens, reinforcing patterns in textual data [33]. Similarly, “Rare Words Heads” prioritize unique or low-frequency tokens [32], while “Previous Heads” capture positional relationships between sequential tokens [25, 12]. Another key attention mechanism, “Retrieval Heads,” boosts the model’s accuracy in locating specific tokens within long passages, a capability often called “Needle-in-a-Haystack” [35]. Together, these examples demonstrate the complex internal structures within LLMs that support a wide range of linguistic functions. However, while much of the focus has been on these attention heads in the context of language, our work shifts the focus to understanding how LLMs process and attend to visual tokens.
@@
Gather in certain layers: Unlike retrieval heads, which appear more universally across layers, visual heads tend to aggregate within specific layers, indicating a unique structural role. This was observed consistently across various model families (LLaMA 2 [31], Phi [1], LLaMA 3 [11], and Mistral [13]), scales (3.8B, 7B, 8B, and 13B).
Concentration: Our initial attempt at analyzing attention weights performed well on a controlled dataset; however, we found that it was inconsistent across different benchmarks. After in-depth analysis with Logic Lens[3], we identified that the concentration, or the sharpness, of the attention heads plays a crucial role in providing a consistent view of the model’s behavior across datasets.
Model-Specific Activation: The location and behavior of visual heads vary across models and training strategies [24], showing that visual heads are not static but adapt dynamically according to training conditions and contexts.
Contextual Responsiveness: Visual heads exhibit dynamic activation based on context, whether visual or textual, enabling them to adjust their function and behavior in response to varying visual and linguistic inputs, improving task performance and adaptability [10].
Enhanced Inference Efficiency: Visual heads offer the potential for more efficient inference, as only a small subset of heads is actively engaged in processing image tokens, which are often dominant in visual tasks. This discovery could lead to optimized models that selectively activate attention heads for visual information, reducing computational load and speeding up inference times. Additionally, our findings explain the effectiveness of certain visual-token compression methods. Specifically, after the initial layers of the LLM, attention becomes highly concentrated on just a few focal areas of the input visual tokens, even if the input consists of thousands of tokens, with the remaining tokens receiving little to no attention.
@@ 2.1 MLLM Architecture
MLLM often consist of modality encoders, an LLM, and an adapter to bridge these components. Modality encoders serve to compress data, such as images or audio, into more concise representations. A widely adopted approach is to leverage pretrained encoders that have been pre-aligned with other modalities. For example, the CLIP [28] model incorporates a visual encoder that has undergone extensive pretraining with text, establishing a semantic alignment between visual and textual modalities. This pre-aligned structure facilitates seamless integration with LLMs. The LLaMA series [30, 31, 11] and the Vicuna family [8] are prominent open-source LLMs that have garnered significant academic interest. Additionally, increasing the parameter size of LLMs can yield further performance gains, much like enhancing input resolution. Liu et al. [20, 19, 21], observe that scaling an LLM from 7B to 13B parameters leads to notable improvements across various benchmarks. Moreover, when scaled up to 34B parameters, the model exhibits emergent zero-shot capability in Chinese, despite being trained solely on English multimodal data.
@@ 2.2 Learnable Connector
This module plays a critical role in bridging the gap between different modalities, facilitating the efficient projection of information into a space that LLMs can interpret. Multimodal information fusion can be broadly categorized into two implementation strategies: token-level and feature-level fusion. In token-level fusion, features output by encoders are transformed into tokens and concatenated with text tokens before entering LLMs. A common and practical approach is to employ a set of learnable query tokens to extract relevant information in a query-driven manner [4], first introduced in BLIP-2 [16] and subsequently adopted by numerous works [7, 9, 38]. This Q-Former-style method condenses visual tokens into a limited number of representation vectors. Conversely, some methods opt for a simpler MLP-based interface to address the modality gap [20, 39, 29, 26]. For instance, the LLaVA series leverages one or two linear MLPs [20, 19] to project visual tokens and align the feature dimensions with a space compatible with LLMs.
@@ 2.3 Attention Head in LLM
Large language models (LLMs) leverage a network of specialized attention heads to enhance their reasoning and interpretative abilities. These attention heads are designed to focus on different aspects of input data, processing information in ways that approximate human cognitive tasks. For instance, the Previous Head is designed to attend to the relationship between current and prior tokens, capturing token sequence patterns that support in-context learning [25]. The Rare Words Head highlights infrequent tokens, aiding models in identifying unique and meaningful details within text [32]. Syntactic Heads identify structural components such as subjects and objects, organizing information by grammatical roles for a more coherent understanding [12]. Additionally, Truthfulness Heads focus on maintaining the factual accuracy of responses, helping the model to produce reliable answers in response to questions [18].
To uncover these attention heads, researchers employ several different methods. Modeling-Free methods, such as zero ablation [27] and activation patching [40] involve modifying latent states and observing their impact on model outputs without training. In contrast, Modeling-Required methods train additional models to assess head functions. These paper primarily emphasizes score-based methods, a Modeling-Free approach that uses mathematical scores, such as Retrieval [35] and Negative [37] Attention Scores, to evaluate specific attention head behaviors directly. This focus on score-based methods provides an efficient framework for understanding specialized functions within LLMs without the need for extensive model adjustments.
@@ 3 Method
Table 1: Summary of model selections, including their LLM family, layer-head architecture, resolution, training strategy, and use of visual tokens.
@@
1.5-7B Vicuna-v1.5 $32\times 32$ $336\times 336$ frozen vision encoder 576
@@
vip-7B $32\times 32$ $336\times 336$ frozen vision encoder 576
@@ 3.1 Experiment Setup
We begin by selecting the specialized dataset, the PointQA dataset [23], mainly for three reasons: (a) Diverse Question Types: PointQA includes three categories of questions, each designed to vary in complexity and focus (as shown in Fig.2): General Questions: These prompt the model to identify objects within specified regions and count instances (e.g., “How many objects are there?”), testing the model’s ability to recognize and count objects. Object-Specific Questions: These directly identify the objects of interest, such as “How many people are in the picture?” This type assesses the model’s ability to detect and count specific objects within a scene. Super Questions: These questions lack explicit object names (e.g., “How many beings are there?”), requiring the model to interpret the scene broadly to infer the correct subject. (b) Region-Specific Questioning and Visual Prompts: PointQA designates specific image regions relevant to each question, which serves two purposes. First, it allows us to quantitatively assess whether the model is focusing on the correct areas, ensuring accurate response generation. Second, by highlighting relevant areas (e.g., with a circle or box), we create visual prompts that guide the model’s attention to specific regions within the image. (c) Controlled Response: By limiting answers to a single numerical token (1-7), PointQA reduces the risk of hallucinated responses, ensuring that answers are directly tied to the visual input rather than influenced by model biases or prior content generation. (d) Look Twice: Consider questions that require spatial reasoning. A natural example is shown on the left in Fig2. The model needs to identify the relevant object in the image first and then use this information to attend to the entire image effectively.
@@ 3.3 Extract Attention Weight
In the PointQA dataset, the target token for each output is a single digit (ranging from 1 to 7). We analyze the attention weights, denoted as $\alpha_{l,h,j}$, which represent the probability that the target token attends to each input token $x_{j}$ for a specific attention head $h$ in layer $l$ of the 7b model. Here, both $l$ and $h$ range from 0 to 32. The attention weights for each head form a $1\times n$ vector, where $n$ is the length of the input sequence plus one (to account for the inclusion of the $\langle s\rangle$ start-of-sequence token). This structure minimizes interference from previously decoded tokens, meaning the target token’s information is derived directly from the input sequence. Unlike previous studies that focus on attention flow [2] or relevance scores [6] to trace connections between input and output tokens, our work emphasizes how individual attention heads attend specifically to image tokens during decoding. This approach provides insights into how attention heads focus on the input image tokens.
@@ 4.1 Attention Weight Analyze
[…]
@@ 4.2 Are Attention Weight Consistent?
We extracted attention weights to four additional popular MLLM benchmarks: MMBench [22], POPE [17], SEED [15], and the recently introduced QBench [34].
As shown in Fig. 5, we observe that while different models exhibit similar attention patterns on Visual General subset, this consistency does not hold for datasets like MMBench and POPE. For these benchmarks, distinctive patterns emerge primarily in the first few layers. To investigate further, we conducted a thorough analysis of the hidden state outputs using Logit Lens [3]. This analysis revealed that in datasets like POPE and MMBench, attention and logit progression stabilize after the initial layers. In contrast, for Visual General, attention weights and logits continue to evolve through the middle layers. This suggests that the answers require more refinement, as the dataset demands “look twice” reasoning, whereas the questions in POPE and MMBench are more straightforward.
@@ 4.3 Concentration
This analysis essentially demonstrates how each attention head localizes its focus across the entire dataset. As shown in the figure, when comparing the plots across rows, we observe that heads with higher concentration or those focusing more on specific parts of the image tend to yield better performance compared to models that do not exhibit this behavior. Furthermore, we found that the 13B family of models displays very similar behavior across the five benchmarks. As shown in Fig. 1, we observe a consistent pattern of model behavior within the dataset. Moreover, we found that integrating concentration with attention weight still holds in our analysis of the distribution of heads across layers as show in Fig7. We can still observe the distribution layer-wise. Additional figures detailing different models and datasets will be included in the supplementary material.
@@ 6 Conclusion
Supplementary Material
7 Logic Lens
To investigate why attention weights are not consistent across different datasets, we employed the “logic lens” to examine the rank of decoded tokens across layers. We aimed to determine if the model still reduces its output to the top-1 guess while comparing other distributions to the final one, specifically by analyzing the rank of the final top-1 guess. This method allows us to see how the model’s predictions evolve from the initial layers to the final output. By examining the rank of decoded tokens layer by layer, we can observe whether the model maintains uncertainty about its predictions in the earlier layers and becomes more certain in the later layers. This granularity helps identify at what stage the model’s predictions start to stabilize.
In the Visual General Benchmark, results become more confirmatory toward the very last layers, indicating that the model in the middle layers is still trying to ascertain the correct answer. However, for the other four benchmarks, the final answer converges to the final result quite rapidly in the initial few layers.This observation explains why, in later layers, the model ceases to distribute attention to image tokens, having already confirmed the answer.
Figure 9: Visualizing Model Behavior Across Datasets: This figure shows different convergence patterns, with blue indicating Rank 1. Particularly in the Visual General Benchmark, the model requires more time to ascertain the correct answer, reflecting a need for Look Twice In contrast, some datasets show a swift stabilization to the final answer, which then persists across layers, showcasing varied strategic adjustments by the model.”
Figure 10: Summation of attention weights for the LLaVA-1.5 Phi-3 Mini 3.8B model, showing a consistent activation in the middle and late layers across various datasets.
Figure 11: Concentration of attention weights sum for the LLaVA-1.5 Phi-3 Mini 3.8B model, illustrating that middle and late layers are predominantly activated across datasets.
8 Masking out Late heads
We observed that, unlike early and middle heads, the later heads have little impact on the final accuracy. Accordingly, we conducted experiments on masking varying numbers of heads across two models, as depicted in the figures below.
Figure 12: This graph displays the performance degradation of v1.5-7b across three different models (MMbench, POPE, Qbench) under two masking strategies (Random and Top) as a function of the percentage of heads masked. The dashed lines represent the random masking strategy, while the solid lines denote the top masking strategy.
From the figures, we observe a consistent trend across both versions 1.5-7b and 1.6-7b: as more heads are masked, there is a decline in accuracy across all three datasets. Notably, even with 100 heads masked, the random masking method still retains relatively high accuracy compared to the baseline model. This finding has significant implications for enhancing the efficiency of key-value (KV) cache and inference processes. Given that image tokens dominate the input (576 out of 650 tokens), it may be feasible to significantly prune the KV cache corresponding to non-essential heads. By computing only for essential heads, we could achieve substantial gains in inference speed.
Furthermore, we noted that the model displays a consistent trend between versions, though the extent of accuracy reduction varies. For example, both MMbench and Qbench exhibit a more noticeable decline in accuracy with Top masking in version 1.6 compared to 1.5, suggesting that version 1.6 is more robust. Additionally, we observed that the rate of decay in accuracy becomes less steep as more of the top heads are masked, indicating that the remaining heads contribute less significantly to the final accuracy.
Figure 13: Similar to Figure 12, this graph presents the impact on v1.6-7b accuracy for the same three models under the same two masking strategies as the masking extent increases. Colored lines differentiate the models, with blue for MMbench, green for POPE, and red for Qbench.
9 Comparative Analysis of Attention
In our analysis, we observed that the 13B model exhibits significant activation predominantly in the middle layers, a pattern that is similarly noted in the 7B model. In contrast, the Phi3 3.8B model, which is the smallest among the models we studied, shows pronounced activation in the later layers. We hypothesize that this distinct behavior in the 3.8B model arises from its limited capacity.
Larger models, such as the 13B and 7B, generally have more parameters and potentially more layers, allowing for a more balanced distribution of the computational load across the network. This distribution often results in substantial activation in the middle layers, where complex feature interactions and transformations occur. In contrast, smaller models like the 3.8B, with fewer parameters and layers, require a more concentrated utilization of available resources. Consequently, the later layers in smaller models may need to undertake more intensive processing tasks compared to their counterparts in larger models.
Figure 14: Concentration times sum of attention weights in the LLaVA v1.5 13B model, indicating that the middle layers are significantly activated across all datasets examined.
Figure 15: Total summation of attention weights across layers in the LLaVA v1.5 13B model. Unlike the smaller 3.8B model, this larger model shows significant activation in middle layers, highlighting a unique pattern where these layers play a more crucial role, particularly in this model configuration.

### v2

@@ 1 Introduction
The integration of large language models (LLMs) into multimodal tasks has led to impressive advancements, with MLLMs, or large vision-language models (LVLMs), achieving notable success in areas ranging from traditional image captioning to more complex visual dialogues. A key element in training these models is the adapter, which connects the visual encoder and the LLM. The standard approach typically begins by freezing both the visual encoder (often based on models like CLIP [clip]) and the LLM, while fine-tuning only the adapter to map visual embeddings into a language-compatible space. Subsequently, the LLM is fine-tuned to follow instructions for visual tasks, enabling it to become a versatile model capable of solving vision problems that traditionally required multiple specialized models [bi2025verifybenchmarkvisualexplanation]. This approach works remarkably well, raising an intriguing question: does the adapter translate visual embeddings into a language space that the LLM can interpret?
@@
Recent studies [fromcliptodino, yang2024law] reveal that representations from encoders like CLIP can be effectively transformed into a space compatible with LLMs by training an adapter, whereas other visual representations face challenges in achieving similar results. This raises a fundamental question about how LLMs, trained exclusively on language embeddings, interact with embeddings originating from fundamentally different visual spaces. In particular, chan2024analyzing emphasizes that the embedding spaces for visual tokens differ significantly from those used for language, as visual tokens do not function as direct counterparts to language tokens but instead occupy a distinct space. Therefore, an in-depth analysis of how LLMs process embeddings from such a divergent space is essential, as current research lacks a clear understanding of how LLMs attend to these visual tokens during inference. To address this problem, it is necessary to explore the attention mechanisms of LLMs when handling visual tokens.
@@
In parallel, research within the LLM domain has increasingly focused on uncovering the internal mechanisms that drive LLM capabilities, particularly around specialized attention heads that contribute distinct functions. For example, “Duplicate Heads” enhance the model’s ability to focus on repeated tokens, reinforcing patterns in textual data [wang2022interpretability]. Similarly, “Rare Words Heads” prioritize unique or low-frequency tokens [voita2019analyzing], while “Previous Heads” capture positional relationships between sequential tokens [olsson2022context, ferrando2024information]. Another key attention mechanism, “Retrieval Heads,” boosts the model’s accuracy in locating specific tokens within long passages, a capability often called “Needle-in-a-Haystack” [wu2024retrieval]. Together, these examples demonstrate the complex internal structures within LLMs that support a wide range of linguistic functions. However, while much of the focus has been on these attention heads in the context of language, our work shifts the focus to understanding how LLMs process and attend to visual tokens.
@@
Gather in certain layers: Unlike retrieval heads, which appear more universally across layers, visual heads tend to aggregate within specific layers, indicating a unique structural role. This was observed consistently across various model families (LLaMA 2 [2023llama2], Phi [abdin2024phi], LLaMA 3 [2024llama3], and Mistral [jiang2023mistral]), scales (3.8B, 7B, 8B, and 13B).
Concentration: Our initial attempt at analyzing attention weights performed well on a controlled dataset; however, we found that it was inconsistent across different benchmarks. After in-depth analysis with Logic Lens[belrose2023eliciting], we identified that the concentration, or the sharpness, of the attention heads plays a crucial role in providing a consistent view of the model’s behavior across datasets.
Model-Specific Activation: The location and behavior of visual heads vary across models and training strategies [naveed2023comprehensive], showing that visual heads are not static but adapt dynamically according to training conditions and contexts.
Contextual Responsiveness: Visual heads exhibit dynamic activation based on context, whether visual or textual, enabling them to adjust their function and behavior in response to varying visual and linguistic inputs, improving task performance and adaptability [de2023visual].
Enhanced Inference Efficiency: Visual heads offer the potential for efficient inference, as only a small subset of heads is actively engaged in processing image tokens, which are often dominant in visual tasks. This discovery could lead to optimized models that activate attention heads for visual information, reducing computational load and speeding up inference times. Additionally, our findings explain the effectiveness of certain visual-token compression methods. Specifically, after the layers of the LLM, attention becomes concentrated on just a few focal areas of the input visual tokens, even if the input consists of thousands of tokens, with the remaining tokens receiving little to no attention.
@@ 2.1 MLLM Architecture
MLLM often consist of modality encoders, an LLM, and an adapter to bridge these components. Modality encoders serve to compress data, such as images or audio, into more concise representations. A widely adopted approach is to leverage pretrained encoders that have been pre-aligned with other modalities. For example, the CLIP [clip] model incorporates a visual encoder that has undergone extensive pretraining with text, establishing a semantic alignment between visual and textual modalities. This pre-aligned structure facilitates seamless integration with LLMs. The LLaMA series [2023llama, 2023llama2, 2024llama3] and the Vicuna family [chiang2023vicuna, Bi_2024] are prominent open-source LLMs that have garnered significant academic interest. Additionally, increasing the parameter size of LLMs can yield further performance gains, much like enhancing input resolution. liu2023llava, liu2023improvedllava, liu2024llavanext, tang2023video, observe that scaling an LLM from 7B to 13B parameters leads to notable improvements across various benchmarks. Moreover, when scaled up to 34B parameters, the model exhibits emergent zero-shot capability in Chinese, despite being trained solely on English data.
@@ 2.2 Learnable Connector
This module plays a critical role in bridging the gap between different modalities, facilitating the efficient projection of information into a space that LLMs can interpret. Multimodal information fusion can be broadly categorized into two implementation strategies: token-level and feature-level fusion. In token-level fusion, features output by encoders are transformed into tokens and concatenated with text tokens before entering LLMs. A common and practical approach is to employ a set of learnable query tokens to extract relevant information in a query-driven manner [carion2020end], first introduced in BLIP-2 [li2023blip] and subsequently adopted by numerous works [x-llm, instructblip, video-llama]. This Q-Former-style method condenses visual tokens into a limited number of representation vectors. Conversely, some methods opt for a simpler MLP-based interface to address the modality gap [liu2023llava, pmc-vqa, pandagpt, detgpt]. For instance, the LLaVA series leverages one or two linear MLPs [liu2023llava, liu2023improvedllava] to project visual tokens and align the feature dimensions with a space compatible with LLMs.
@@ 2.3 Attention Head in LLM
Large language models (LLMs) leverage a network of specialized attention heads to enhance their reasoning and interpretative abilities. These attention heads are designed to focus on different aspects of input data, processing information in ways that approximate human cognitive tasks. For instance, the Previous Head is designed to attend to the relationship between current and prior tokens, capturing token sequence patterns that support in-context learning [olsson2022context]. The Rare Words Head highlights infrequent tokens, aiding models in identifying unique and meaningful details within text [voita2019analyzing]. Syntactic Heads identify structural components such as subjects and objects, organizing information by grammatical roles for a more coherent understanding [ferrando2024information]. Additionally, Truthfulness Heads focus on maintaining the factual accuracy of responses, helping the model to produce reliable answers in response to questions [liang2024internal].
To uncover these attention heads, researchers employ several different methods. Modeling-Free methods, such as zero ablation [pochinkov2024investigating] and activation patching [zheng2024attention] involve modifying latent states and observing their impact on model outputs without training. In contrast, Modeling-Required methods train additional models to assess head functions. These paper primarily emphasizes score-based methods, a Modeling-Free approach that uses mathematical scores, such as Retrieval [wu2024retrieval] and Negative [yu2024correcting] Attention Scores, to evaluate specific attention head behaviors directly. This focus on score-based methods provides an efficient framework for understanding specialized functions within LLMs without the need for extensive model adjustments.
@@ 3 Method
Table 1: Summary of models, including their LLM family, layer-head architecture, resolution, training strategy, and use of visual tokens.
@@
1.5-7B Vicuna-v1.5 $2\times 32 33$$36\times 336$ frozen vision encoder 576
@@
vip-7B $2\times 32 33$$36\times 336$ frozen vision encoder 576
@@ 3.1 Experiment Setup
We begin by selecting the specialized dataset, the PointQA dataset [mani2020point], mainly for three reasons: (a) Diverse Question Types: PointQA includes three categories of questions, each designed to vary in complexity and focus (as shown in Fig.2): General Questions: These prompt the model to identify objects within specified regions and count instances (e.g., “How many objects are there?”), testing the model’s ability to recognize and count objects. Object-Specific Questions: These directly identify the objects of interest, such as “How many people are in the picture?” This type assesses the model’s ability to detect and count specific objects within a scene. Super Questions: These questions lack explicit object names (e.g., “How many beings are there?”), requiring the model to interpret the scene broadly to infer the correct subject. (b) Region-Specific Questioning and Visual Prompts: PointQA designates specific image regions relevant to each question, which serves two purposes. First, it allows us to quantitatively assess whether the model is focusing on the correct areas, ensuring accurate response generation. Second, by highlighting relevant areas (e.g., with a circle or box), we create visual prompts that guide the model’s attention to specific regions within the image. (c) Controlled Response: By limiting answers to a single numerical token (1-7), PointQA reduces the risk of hallucinated responses, ensuring that answers are directly tied to the visual input rather than influenced by model biases or prior content generation. (d) Look Twice: Consider questions that require spatial reasoning. A natural example is shown on the left in Fig2. The model needs to identify the relevant object in the image first and then use this information to attend to the entire image effectively.
@@ 3.3 Extract Attention Weight
In the PointQA dataset, the target token for each output is a single digit (ranging from 1 to 7). We analyze the attention weights, denoted as $\alpha_{l,h,j}$, which represent the probability that the target token attends to each input token $x_{j}$ for a specific attention head $h$ in layer $l$ of the 7b model. Here, both $l$ and $h$ range from 0 to 32. The attention weights for each head form a $1\times n$ vector, where $n$ is the length of the input sequence plus one (to account for the inclusion of the $\langle s\rangle$ start-of-sequence token). This structure minimizes interference from previously decoded tokens, meaning the target token’s information is derived directly from the input sequence. Unlike previous studies that focus on attention flow [Abnar_2020] or relevance scores [Chefer_2021] to trace connections between input and output tokens, our work emphasizes how individual attention heads attend specifically to image tokens during decoding. This approach provides insights into how attention heads focus on the input image tokens.
@@ 4.1 Attention Weight Analyze
For each sample, we recorded attention weights across layers $l$, heads $h$, yielding an array of dimensions $l\times h\times\text{seq\_len}$, where seq_len has an average length of 650 tokens in the PointQA dataset, including 576 tokens representing image regions. This allows us to partition attention analysis into two primary regions: the image token region and the question region token, which refers to the image token inside the bounding box from the question. For each sample, we computed two key metrics: i The total attention weight in the image region. ii The total attention weight in the question bounding box (bbox). We then averaged these values across all 5,700 samples to obtain a comprehensive view of attention patterns across different model configurations, as visualized in Figure 3. We observe several distinct patterns from the image.
@@ 4.2 Are Attention Weight Consistent?
We extracted attention weights to four additional popular MLLM benchmarks: MMBench [liu2025mmbench], POPE [li2023evaluating], SEED [li2023seed], and the recently introduced QBench [wu2023q].
As shown in Fig. 5, we observe that while different models exhibit similar attention patterns on Visual General subset, this consistency does not hold for datasets like MMBench and POPE. For these benchmarks, distinctive patterns emerge primarily in the first few layers. To investigate further, we conducted a thorough analysis of the hidden state outputs using Logit Lens [belrose2023eliciting]. This analysis revealed that in datasets like POPE and MMBench, attention and logit progression stabilize after the initial layers. In contrast, for Visual General, attention weights and logits continue to evolve through the middle layers. This suggests that the answers require more refinement, as the dataset demands “look twice” reasoning, whereas the questions in POPE and MMBench are more straightforward.
@@ 4.3 Concentration
This analysis essentially demonstrates how each attention head localizes its focus across the entire dataset. As shown in the figure, when comparing the plots across rows, we observe that heads with higher concentration or those focusing more on specific parts of the image tend to yield better performance compared to models that do not exhibit this behavior. Furthermore, we found that the 13B family of models displays very similar behavior across the five benchmarks. As shown in Fig. 1, we observe a consistent pattern of model behavior within the dataset. Moreover, we found that integrating concentration with attention weight still holds in our analysis of the distribution of heads across layers as show in Fig7. We can still observe the distribution layer-wise. Additional figures detailing different models and datasets will be included in the supplementary.
@@ 6 Conclusion
7 Acknowledgment
This work was supported in part by the Defense Advance Research Projects Agency under contract number HR00112220003. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies; no official endorsement should be inferred.